# Towards SLO-Optimized LLM Serving via Automatic Inference Engine Tuning

### v1

@@ IV-A SCOOT Workflow
The workflow of SCOOT is depicted in Fig. 3, involving nine key steps to find optimized parameter configurations. Customers define their optimization objectives and SCOOT ❶ gathers their request traces that include both input text of requests and output text generated by the LLM. Then, SCOOT leverages Sobol sequence-based Quasi-Monte Carlo [19] to uniformly ❷ sample configurations across the search space, where the number of samples matches the search space’s dimensionality. Then, SCOOT runs the inference engine with each sampled configuration and stress tests the inference service to obtain initial observations. These observations of configuration-performance pairs are leveraged to ❸ build a random forest and ❹ construct a surrogate model to learn the $POF(\cdot)$ and predict the probability distribution of each optimization objective, respectively. Subsequently, acquisition functions are exploited to ❺ assess configurations according to the predicted results. Based on the assessment, a solver ❻ suggests multiple configurations in parallel while adhering to the known constraints and ❼ ensuring hidden constraints using $POF(\cdot)$ learned by the random forest. Lastly, the inference engine ❽ is started with each suggested parameter configuration, and the stress testing is conducted to ❾ obtain new observations to refine the random forest and the surrogate model. Steps ❸$\sim$ run iteratively, and the tuning process stops until the number of observations reaches a given threshold.
@@
In the above workflow, the LLM, GPU number, and GPU type used for tuning are the same as the LLM inference service owned by the customer.
IV-B Bayesian Optimization-based Solution
@@
BO is a theoretically grounded method for solving the problem of maximizing black-box objective functions. BO leverages a probabilistic surrogate model to approximate the objective function and use acquisition functions to access configuration points for suggestion. BO can avoid excessive evaluations by exploring the complex multi-dimensional search space efficiently and intelligently [13], which is suitable for solving problems with expensive evaluation overhead.
@@
SCOOT leverages SOBO to minimize tail latency for non-interactive online applications and maximize throughput for periodically invoked applications, respectively. For interactive applications, SCOOT exploits MOBO to minimize TTFT and TPOT simultaneously. MOBO aims to find a set of parameter configurations representing the Pareto frontier that denotes the optimal tradeoffs between TTFT and TPOT.
(a) TTFT-only optimization.
(b) TPOT-only optimization.
Figure 4: Contradictory TTFT and TOPT. Decreased TTFT often leads to increased TPOT, and vice versa. The red line indicates the objective value under the default parameter configuration.
We do not linearly combine TTFT and TPOT as a single optimization objective and solve it with SOBO for the following three reasons. First, the scale of TTFT and TPOT are significantly different. TTFT is hundreds or even thousands of times larger than TPOT. Second, it is difficult to assign weights to TTFT and TPOT to make a good trade-off. Third, under stress testing, TTFT and TPOT are always two conflicting optimization objectives.
We separately take TTFT and TPOT as the only optimization objectives and conduct performance tuning for inference services of LLAMA2-13B deployed on various computing resources using the BOT request trace. As shown in Fig. 4, the experimental results confirm that for TTFT and TPOT, optimizing for one always compromised the other. For request traces of other applications, we also observe similar results.
The reason is that in order to decrease TPOT, it is significant to decrease the batch size controlled by the parameter max-num-seqs to reduce the per-token latency, which greatly hurts the request throughput and causes a long queuing time for requests, thus leading to a large TTFT. Conversely, to reduce TTFT, a large batch size is always set to allow newly arrived requests to be processed promptly without queuing for a long time. However, a large batch size often slows down inference, leading to increased TPOT.
Therefore, for interactive online applications, optimizing the Pareto frontier of TPOT and TTFT is appropriate. Besides, this can offer customers a wide range of configurations to choose from, providing a high flexibility of SLO optimization.
IV-C Surrogate Model
@@
The surrogate model of BO is responsible for predicting the objective function $f(\boldsymbol{x})$ on the basis of observations. It models $f(\boldsymbol{x})$ for a given configuration $\boldsymbol{x}$ as a random variable and predicts its probability distribution. In the context of performance tuning for LLM inference engines, $f(\boldsymbol{x})$ can represent the objective functions of request throughput $T(\boldsymbol{x})$, request tail latency $L(\boldsymbol{x})$, average TTFT $\Phi(\boldsymbol{x})$, and average TPOT $\Theta(\boldsymbol{x})$.
SCOOT uses the Gaussian process (GP) as the surrogate model because it is parameter-free and provides closed-form calculation. GP is characterized by a mean function $m(\boldsymbol{x})$ and a covariance function $k(\boldsymbol{x},\boldsymbol{x^{\prime}})$. Formally, a GP is defined as
$f(\boldsymbol{x})\sim GP(m(\boldsymbol{x}),k(\boldsymbol{x},\boldsymbol{x^{\prime}})),$ (4)
where the mean function $m(\boldsymbol{x})$ corresponds to the expected value of $f(\boldsymbol{x})$ for a specific input $\boldsymbol{x}$, and the covariance function $k(\boldsymbol{x},\boldsymbol{x^{\prime}})$ models the similarity between objective values at different inputs $\boldsymbol{x}$ and $\boldsymbol{x^{\prime}}$. $k(\boldsymbol{x},\boldsymbol{x^{\prime}})$ can capture the correlation structure of the objective function and control the variability and smoothness of the GP. SCOOT employs the Matern 3/2 kernel function with input wrapping as the covariance function due to its powerful capability to balance the smoothness and flexibility when modeling unknown functions [20, 21].
@@
Given a parameter configuration $\boldsymbol{x}$, GP assumes that the probability distribution of $f(\boldsymbol{x})$ follows a Gaussian distribution whose mean $\mu(f(\boldsymbol{x}))$ is calculated by
$\mu(f(\boldsymbol{x}))=m(\boldsymbol{x})+k(\boldsymbol{x},\boldsymbol{\mathcal{X}})(k(\boldsymbol{\mathcal{X}},\boldsymbol{\mathcal{X}})+\tau^{2}I)^{-1}(\boldsymbol{\mathcal{Y}}-m(\boldsymbol{\mathcal{X}})),$ (5)
and the variance $\sigma^{2}(f(x))$ is computed by
$\sigma^{2}(f(\boldsymbol{x}))=k(\boldsymbol{x},\boldsymbol{x})-k(\boldsymbol{x},\boldsymbol{\mathcal{X}})(k(\boldsymbol{\mathcal{X}},\boldsymbol{\mathcal{X}})+\tau^{2}I)^{-1}k(\boldsymbol{\mathcal{X}},\boldsymbol{x}),$ (6)
where $\tau^{2}$ is the level of white noise, $\boldsymbol{\mathcal{X}}$ represents the previously evaluated, configurations and $\boldsymbol{\mathcal{Y}}$ denotes the corresponding observed objective values.
@@
IV-D Acquisition Function
The acquisition function of BO assesses parameter configurations in the search space based on the predicted results provided by the surrogate model. The acquisition function calculates a score for a configuration point $\boldsymbol{x}$ according to the GP’s predicted mean $\mu(f(\boldsymbol{x}))$ (indicating the expected performance) and predicted variance $\sigma^{2}(f(\boldsymbol{x}))$ (representing uncertainty). The parameter configurations with high scores are more likely to be suggested for the next evaluation. Different acquisition functions balance exploration (visiting areas with high uncertainty) and exploitation (intensively searching areas with good known objective values) in different manners.
IV-D1 SOBO Acquisition Function
For SOBO, commonly used acquisition functions include upper confidence bound (UCB), probability of improvement (PI), and expected improvement (EI). UCB incorporates both mean and variance into the score and prioritizes the point with a high balance of exploration and exploitation through a tunable trade-off parameter $\beta$, which is expressed by
$UCB(\boldsymbol{x})=\mu(f(\boldsymbol{x}))+\beta\cdot\sigma(f(\boldsymbol{x})),$ (7)
where $\sigma(f(\boldsymbol{x}))$ represents the standard deviation and it is calculated by $\sqrt{\sigma^{2}(f(\boldsymbol{x}))}$.
PI prioritizes the configuration point that is likely to yield an improvement over the current known best observation, which is expressed by
$PI(\boldsymbol{x})=P(f(\boldsymbol{x})+\xi>f(\boldsymbol{x}^{+})),$ (8)
where $\boldsymbol{x}^{+}$ is the configuration point that has the largest objective function value known so far, and $\xi$ is a parameter leveraged to encourage exploration.
@@ IV-D1 SOBO Acquisition Function
$EI(\boldsymbol{x})=\mathbb{E}(\max(0,f(\boldsymbol{x})+\xi-f(\boldsymbol{x}^{+}))),$ (9)
where $\boldsymbol{x}^{+}$ is the best observed configuration point, and $\xi$ is also used to encourage exploration.
@@
Figure 5: An illustration of two-dimensional HV. $f_{1}$ and $f_{2}$ are two objective functions and $\boldsymbol{r}$ is the reference point. (a) The blue area represents the HV of the existing solution set. (b) The yellow area depicts the HV improvement after adding $\boldsymbol{y}_{4}$.
IV-D2 MOBO Acquisition Function
For MOBO, expected hypervolume improvement (EHVI) is widely adopted as the acquisition function to identify promising parameter configurations that balance the conflicting objectives. EHVI assesses a new configuration’s quality based on the potential improvement it might bring to the hypervolume of the existing solution set, and EHVI is calculated by
$EHVI(\boldsymbol{x})=\mathbb{E}(\max(0,HV(\boldsymbol{\mathcal{Y}}\cup\{f(\boldsymbol{x})\})-HV(\boldsymbol{\mathcal{Y}}))),$ (10)
@@ IV-D2 MOBO Acquisition Function
The hypervolume function $HV(\cdot)$ is calculated by measuring the volume of space enclosed between the Pareto frontier and a reference point $\boldsymbol{r}$, where $\boldsymbol{r}$ is typically a lower bound point for all objective functions. When there are only two optimization objectives, $HV(\cdot)$ is calculated by summing the areas of rectangles formed between each point on the Pareto frontier and the selected reference point $\boldsymbol{r}$. Figure 5 presents how the two-dimensional $HV(\cdot)$ is computed.
@@
A larger $HV(\cdot)$ value indicates a superior Pareto front as it covers a larger objective space, offering a broader range of promising solutions. Therefore, by maximizing the EVHI, the search process can be guided towards configurations that can improve the overall quality and diversity of the Pareto frontier, thus optimizing multiple objectives simultaneously and providing more excellent optional solutions.
IV-E Configuration Suggestion
IV-E1 SOBO Suggestion
@@
For SOBO, a common practice is to select a specific acquisition function and suggest parameter configurations that maximize the acquisition function using a solver. However, when it comes to real-world applications, the selected acquisition function might work sub-optimally for the tuning task, and the best-performing acquisition function is challenging to identify in advance [22].
@@
To mitigate this issue, SCOOT employs multi-objective acquisition function ensemble (MACE) [23]. MACE uses three acquisition functions, UCB, PI, and EI, to assess configuration points and runs a solver to find out the Pareto frontier that denotes the optimal trade-offs of the three acquisition functions. Suggested configurations are randomly selected in the Pareto frontier. With MACE, SCOOT can avoid using sub-optimal acquisition functions all the time, thereby improving the quality of suggested parameter configurations.
@@
During the tuning process, for UCB, $\beta$ is dynamically adjusted, and in the $t$th tuning iteration, $\beta$ is calculated by ${2log(\frac{t^{2}\pi^{2}}{6\delta})}$, where $\delta=\frac{1}{t^{2}}$ [24]. For EI and PI, $\xi$ is set to a fixed value of 0.0001.
IV-E2 MOBO Suggestion
@@
For MOBO, SCOOT leverages the EHVI as the only acquisition function and runs a solver to suggest parameter configurations by maximizing EHVI, where the default parameter configuration of the inference engine is chosen as the reference point $\boldsymbol{r}$.
IV-E3 Known Constraint Assurance
@@
For both SOBO and MOBO, in the process of solvers solving optimization problems, the solution space is pruned by known constraints to avoid engine startup failure, and hence the known constraints can be guaranteed.
Input: $N$: the total number of observations;
$k$: parallelism degree;
@@ 2 $\mathcal{O}\leftarrow$ UNIFORM_SAMPLE_AND_EVALUATE()
// Train random forest regression
$RF\leftarrow$ TRAIN($\mathcal{O}$) // Train Random Forest
@@
            $c$ $\leftarrow 0$ /* Clear continuous feasible suggestion count */
@@ 15                  $c\leftarrow c-5$
      $RF\leftarrow$ TRAIN($\mathcal{O}$) // Refine RF
@@
IV-F Parallel Configuration Suggestion
With the popularity of LLMs, an increasing number of LLM inference services are being deployed, and tuning them incurs substantial time overhead. To mitigate this problem, SCOOT suggests multiple parameter configurations at a time and utilizes multiple sets of computing resources to conduct configuration evaluations in parallel.
Given the parallelism degree $k$, for SOBO, SCOOT randomly selects $k$ points from the Pareto frontier of the three acquisition functions for suggesting. For MOBO, the top-$k$ points with the largest EHVI are chosen for suggesting. Thus, tuning can be accelerated by a factor of $k$ when the number of observations is fixed.
Although parallel suggestions require additional computing resources to evaluate configurations in parallel, many computing resources in production clusters are often idle during off-peak hours such as nights and weekends, which can be conveniently used for tuning the inference engine.
IV-G Random Forest-based POF Learning
Frequent violations of hidden constraints can result in a significant number of invalid observations caused by engine crashes, which not only hurts the tuning efficiency but also hampers the surrogate model from adequately learning the relationship between the objective functions and configuration points, thereby hurting the optimization of SLOs. To handle hidden constraints, a common practice is to assign penalty objective values to the parameter configurations that cause engine crashes. However, it is challenging to set appropriate penalty objectives and such penalty objective values often negatively affect the surrogate model [25]. Therefore, SCOOT leverages random forest regression to learn the $POF(\cdot)$ for parameter configurations and handle hidden constraints by limiting the probability of feasibility to be greater than or equal to a given threshold, which is expressed by Eq. (3) in Section III. By restricting the solver from adhering to the hidden constraints during the configuration suggestion, SCOOT can substantially reduce the number of invalid observations without affecting the surrogate model.
For the hidden constraint expressed in Eq. (3), the feasibility probability threshold $\Delta$ is critical. If $\Delta$ is fixed, setting it too high may result in repeatedly suggesting configuration points near feasible configurations, potentially missing out on other superior configuration points. Conversely, setting the threshold too low may lead to suggestions of many infeasible configurations. Therefore, SCOOT dynamically adjusts the threshold $\Delta$ during the tuning process.

### v2

@@ 4.1. SCOOT Workflow
The workflow of SCOOT is depicted in Fig. 3, which consists of nine key steps to find optimized parameter configurations. Customers define their optimization objectives, and SCOOT ❶ gathers their request traces that include both input text of requests and output text generated by the LLM. Then, SCOOT leverages Sobol sequence-based Quasi-Monte Carlo (caflisch1998monte, ) to uniformly ❷ sample configurations across the search space, where the number of samples matches the search space’s dimensionality. Then, SCOOT runs the inference engine with each sampled configuration and stress tests the inference service to obtain initial observations. These observations of configuration-performance pairs are leveraged to ❸ build a random forest and ❹ construct a surrogate model to learn the $POF(\cdot)$ and predict the probability distribution of each optimization objective, respectively. Subsequently, acquisition functions are exploited to ❺ assess configurations according to the predicted results. Based on the assessment, a solver ❻ suggests multiple configurations in parallel while adhering to the known constraints and ❼ ensuring hidden constraints using $POF(\cdot)$ learned by the random forest. Lastly, the inference engine ❽ is started with each suggested parameter configuration, and the stress testing is conducted to ❾ obtain new observations to refine the random forest and the surrogate model. Steps ❸$\sim$ run iteratively, and the tuning process stops until the number of observations reaches a given threshold.
@@
In the workflow, the LLM, GPU number, and GPU type used for tuning are the same as the inference service owned by the customer.
4.2. Bayesian Optimization-based Solution
@@
BO is a theoretically grounded method for finding the optimum of black-box functions. It can explore the complex multi-dimensional search space efficiently and intelligently (wang2023recent, ), which is suitable for solving problems with expensive evaluation overhead. BO leverages a surrogate model to approximate the objective function and use acquisition functions to assess configuration points for suggesting.
@@
SCOOT leverages SOBO to maximize throughput and minimize tail latency for non-interactive offline and online applications, respectively. For interactive applications, it exploits MOBO to minimize TTFT and TPOT simultaneously by finding a set of parameter configurations representing the Pareto frontier that denotes the optimal trade-offs between TTFT and TPOT. We do not linearly combine TTFT and TPOT as a single optimization objective and solve it with SOBO because TTFT can be hundreds or even thousands of times larger than TPOT. Hence, it is difficult to assign weights to TTFT and TPOT to make a good trade-off. Besides, TTFT and TPOT are always two conflicting optimization objectives as illustrated in Appendix A.1.
4.3. Surrogate Model
@@
The surrogate model of BO predicts the objective function $f(\boldsymbol{x})$ based on observations. It models $f(\boldsymbol{x})$ for a given configuration $\boldsymbol{x}$ as a random variable and predicts its probability distribution. In the context of LLM inference engine tuning, $f(\boldsymbol{x})$ can represent the objective functions of request throughput $T(\boldsymbol{x})$, request tail latency $L(\boldsymbol{x})$, average TTFT $\Phi(\boldsymbol{x})$, and average TPOT $\Theta(\boldsymbol{x})$.
@@
SCOOT uses the Gaussian process (GP) as the surrogate model. Given a parameter configuration $\boldsymbol{x}$, GP assumes that the probability distribution of $f(\boldsymbol{x})$ follows a Gaussian distribution whose mean $\mu(f(\boldsymbol{x}))$ and the variance $\sigma^{2}(f(\boldsymbol{x}))$ are respectively computed by
(4) $\mu(f(\boldsymbol{x}))=k(\boldsymbol{x},\boldsymbol{\mathcal{X}})(k(\boldsymbol{\mathcal{X}},\boldsymbol{\mathcal{X}})+\tau^{2}I)^{-1}\boldsymbol{\mathcal{Y}},$
(5) $\sigma^{2}(f(\boldsymbol{x}))=k(\boldsymbol{x},\boldsymbol{x})-k(\boldsymbol{x},\boldsymbol{\mathcal{X}})(k(\boldsymbol{\mathcal{X}},\boldsymbol{\mathcal{X}})+\tau^{2}I)^{-1}k(\boldsymbol{\mathcal{X}},\boldsymbol{x}),$
where $\boldsymbol{\mathcal{X}}$ represents the previously evaluated configurations, $\boldsymbol{\mathcal{Y}}$ denotes the corresponding observed objective values, $\tau^{2}$ is the level of white noise, and $k(\boldsymbol{x},\boldsymbol{x^{\prime}})$ is the covariance function that quantifies the similarity between input points $\boldsymbol{x}$ and $\boldsymbol{x^{\prime}}$ for inferring the relationships between their objective function values. SCOOT employs the Matern kernel ($\frac{3}{2}$) with input wrapping (snoek2014input, ) as the covariance function due to its capability to balance the smoothness and flexibility when modeling unknown functions. SCOOT utilizes maximum likelihood estimation (seeger2004gaussian, ) to learn $\tau^{2}$ during tuning.
@@
4.4. Acquisition Function
The acquisition function of BO assesses parameter configurations in the search space, which calculates a score for each configuration point $\boldsymbol{x}$ according to the surrogate model’s predicted mean $\mu(f(\boldsymbol{x}))$ (indicating the expected performance) and variance $\sigma^{2}(f(\boldsymbol{x}))$ (representing uncertainty). Parameter configurations with high scores are more likely to be suggested for evaluation. Different acquisition functions balance exploration (visiting areas with high uncertainty) and exploitation (intensively searching areas with good known objective values) in different manners.
4.4.1. SOBO Acquisition Function
For SOBO, commonly used acquisition functions include upper confidence bound (UCB), probability of improvement (PI), and expected improvement (EI). UCB incorporates both mean and variance into the score and prioritizes the point with a high balance of exploration and exploitation through a trade-off parameter $\beta$, which is expressed by
(6) $UCB(\boldsymbol{x})=\mu(f(\boldsymbol{x}))+\beta\cdot\sigma(f(\boldsymbol{x})),$
where $\sigma(f(\boldsymbol{x}))=\sqrt{\sigma^{2}(f(\boldsymbol{x}))}$ represents the standard deviation.
PI prioritizes the point that is likely to yield an improvement over the current known best observation, which is expressed by
(7) $PI(\boldsymbol{x})=P(f(\boldsymbol{x})+\xi>f(\boldsymbol{x}^{+})),$
where $\boldsymbol{x}^{+}$ is the point that has the largest objective function value known so far, and $\xi$ is leveraged to encourage exploration.
@@ 4.4.1. SOBO Acquisition Function
(8) $EI(\boldsymbol{x})=\mathbb{E}(\max(0,f(\boldsymbol{x})+\xi-f(\boldsymbol{x}^{+}))),$
where $\boldsymbol{x}^{+}$ represents the best observed configuration point as well, and $\xi$ is also a parameter used to encourage exploration.
@@
Figure 4. Illustration of two-dimensional HV. $f_{1}$ and $f_{2}$ are two objective functions and $\boldsymbol{r}$ is the reference point. (a) Blue area represents the HV of the existing solution set. (b) Yellow area depicts the HV improvement after adding $\boldsymbol{y}_{4}$.
4.4.2. MOBO Acquisition Function
For MOBO, expected hypervolume improvement (EHVI) is widely adopted as the acquisition function to identify promising parameter configurations that denote optimal trade-offs between multiple optimization objectives. EHVI assesses a new configuration point based on the potential improvement it might bring to the hypervolume of the existing solution set, which is calculated by
(9) $EHVI(\boldsymbol{x})=\mathbb{E}(\max(0,HV(\boldsymbol{\mathcal{Y}}\cup\{f(\boldsymbol{x})\})-HV(\boldsymbol{\mathcal{Y}}))),$
@@ 4.4.2. MOBO Acquisition Function
The hypervolume function $HV(\cdot)$ is calculated by measuring the volume of space enclosed between the Pareto frontier and a reference point $\boldsymbol{r}$, where $\boldsymbol{r}$ is typically a lower bound point for all objective functions. When there are only two optimization objectives, $HV(\cdot)$ is calculated by summing the areas of rectangles formed between each point on the Pareto frontier and $\boldsymbol{r}$. Figure 4 presents how the two-dimensional $HV(\cdot)$ is computed.
@@
A larger $HV(\cdot)$ value indicates a Pareto front that covers a larger objective space, offering a broader range of promising solutions. By maximizing the EVHI, the search process can be guided towards configurations that can improve the overall quality and diversity of the Pareto frontier, thus optimizing multiple objectives simultaneously and providing more optional configurations for customers.
4.5. Configuration Suggestion
4.5.1. SOBO Suggestion
@@
For SOBO, a common practice is to select a specific acquisition function and suggest parameter configurations that maximize the acquisition function. However, when it comes to real-world applications, the selected acquisition function might work sub-optimally for the tuning task, and the best acquisition function is challenging to identify beforehand (cowen2022hebo, ).
@@
To mitigate this issue, SCOOT employs multi-objective acquisition function ensemble (MACE) (zhang2021efficient, ). MACE uses three acquisition functions, UCB, PI, and EI, to assess configuration points and runs a solver to find out the Pareto frontier that denotes the optimal trade-offs of the three acquisition functions. Suggested configurations are randomly selected in the Pareto frontier. With MACE, SCOOT can avoid using sub-optimal acquisition functions all the time, thereby improving the quality of suggested parameter configurations.
@@
During the tuning process, for UCB, $\beta$ is dynamically adjusted, and in the $t$th tuning iteration, $\beta$ is calculated by ${2log(\frac{t^{2}\pi^{2}}{6\delta})}$, where $\delta=\frac{1}{t^{2}}$ (srinivas2010gaussian, ). For EI and PI, $\xi$ is set to a fixed value of 0.0001.
4.5.2. MOBO Suggestion
@@
For MOBO, SCOOT leverages the EHVI as the acquisition function and runs a solver to suggest parameter configurations by maximizing EHVI. The default configuration of the inference engine is chosen as the reference point $\boldsymbol{r}$.
4.5.3. Known Constraint Assurance
@@
For both SOBO and MOBO, in the process of solvers solving optimization problems, the solution space is pruned by known constraints.
4.6. Parallel Configuration Suggestion
With the popularity of LLMs, an increasing number of LLM inference services are being deployed, and tuning them incurs substantial time overhead. To mitigate this problem, SCOOT suggests multiple parameter configurations at a time and utilizes multiple sets of computing resources to conduct stress testing in parallel.
Given the parallelism degree $k$, for SOBO, SCOOT randomly selects $k$ points from the Pareto frontier of the three acquisition functions for suggesting. For MOBO, the top-$k$ points with the largest EHVI are suggested. Thus, tuning can be accelerated by a factor of $k$ when the total observation number is fixed.
Although the parallel suggestion requires additional computing resources to evaluate configurations in parallel, many computing resources in production clusters are often idle during off-peak hours such as nights and weekends, which can be conveniently used for tuning LLM inference engines.
Input: $N$: total observation number; $k$: parallelism degree;
@@ 2 $\mathcal{O}\leftarrow$ UNIFORM_SAMPLE_AND_EVALUATE()
$RF\leftarrow$ TRAIN($\mathcal{O}$) // Train random forest regression
@@
            $c$ $\leftarrow 0$ // Clear feasible suggestion count
@@ 15                  $c\leftarrow c-5$
      $RF\leftarrow$ TRAIN($\mathcal{O}$) // Refine Random Forest
@@
4.7. Random Forest-based POF Learning
To handle hidden constraints, a common practice is to assign penalty objective values to the infeasible parameter configurations that cause engine crashes. However, it is challenging to set appropriate penalty objective values. Besides, penalty objective values often hurt the surrogate model (hellsten2023baco, ). Therefore, SCOOT leverages random forest regression to learn the $POF(\cdot)$ and handle hidden constraints by restricting the solver to adhere to Eq. (3) in the process of configuration suggestion. In this way, SCOOT can substantially reduce invalid observations without affecting the surrogate model.
For the hidden constraint expressed in Eq. (3), the feasibility probability threshold $\Delta$ is critical. If $\Delta$ is fixed, setting it too high may result in repeatedly suggesting configuration points near already observed feasible configurations, potentially missing out on other superior configuration points. Conversely, setting $\Delta$ too low may lead to suggestions of infeasible configurations. Hence, SCOOT dynamically adjusts the threshold $\Delta$ during the tuning process.